OpenStudy (anonymous):

What time will it be 7 hours 53 minutes after 19:20?

sammixboo (sammixboo):

Well, how much more time it is until it is 24:00 from 19:20?

OpenStudy (kendricklamar2014):

27:13

OpenStudy (anonymous):

3:13

sammixboo (sammixboo):

There is no such time as 27:13 :P

sammixboo (sammixboo):

There is only 24 hours in a day, not 27

OpenStudy (kendricklamar2014):

oops my bad

OpenStudy (kendricklamar2014):

Yes it will be 3:13 like @minimissy22 said

OpenStudy (aum):

7 hours 53 minutes after 19:20? Add the hours and minutes separately: 19 + 7 = 26 hours minus 24 hours = 2:00 53 + 20 = 73 minutes = 1hour + 13 minutes = 1:13 Combine them: 2:00 + 1:13 = 3:13